OpenStudy (adrianna.gongora):

Regan is trying to find the equation of a quadratic that has a focus of (–2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (adrianna.gongora):

f(x) = 3(x + 4)^2 + 1 g(x) = 2x^2 - 16x + 15 h(x) = a(x - 1) - 3 axis of symmetry? Axis of symmetry for f(x) is x=-4 For g(x) Taking the derivative g'(x) = 4x - 16 = 4(x - 4) Axis of symmetry for g(x) is x=4 Axis of symmetry for h(x) is x=1 Is the answer -.- I figured it out.